# QMDB: Quick Merkle Database

### v1

@@
QMDB: Quick Merkle Database
Isaac Zhang     Ryan Zarick     Daniel Wong     Thomas Kim     Bryan Pellegrino
Mignon Li     Kelvin Wong
LayerZero Labs
@@ 1 Introduction
Updating, managing, and proving world state are key bottlenecks facing the execution layer in modern blockchains. Within the execution layer, the storage layer, in particular, has traditionally traded off performance (throughput) and decentralization (capital and infrastructure barriers to participation). Blockchains typically implement active state management using an Authenticated Data Structure (ADS) such as a Merkle Patricia Trie (MPT). Unfortunately, typical MPT-based ADSes incur a high amount of write amplification (WA) with many costly random writes for each state update, which requires storing the entire structure in DRAM to avoid getting bottlenecked by the SSD. As a result, the performance and scaling of blockchains is I/O-bound, and the key to unlocking higher performance with larger datasets is to optimize the use of SSD IOPS more efficiently by reducing WA.
We present Quick Merkle Database (QMDB), a resource efficient SSD-optimized ADS with in-memory Merkleization that implements a superset of the app-level features of existing RocksDB-backed MPT ADSes with 6×6$\times$6 × throughput on large datasets. QMDB implements state reads with 1 SSD read, state updates with O(1) SSD IO, and in-memory Merkleization (0 SSD reads or writes), all on a DRAM footprint small enough to run on consumer-grade PCs. Notably, it has a constant storage complexity (in terms of DRAM).
@@
QMDB eliminates this inefficiency by unifying the world state and Merkle tree storage, persisting all state updates in an append-only log, and eliminating all SSD reads and writes from Merkleization. By grouping updates into fixed-size immutable subtrees called twigs, QMDB can Merkleize state updates without reading or writing any world state; this essentially compresses the Merkle tree by several orders of magnitude, allowing it to be stored in a modest amount of DRAM.
@@
Figure 1: Entries are inserted sequentially into the Fresh twig, which eventually transitions into the Full state. As Entries are deleted, Full twigs become Inactive, then transition to Pruned. Upper nodes are recursively pruned after both of their children are pruned.
@@ 2 Background
MPTs combine the efficient proof generation of the Merkle tree with the fast lookups of the Patricia trie and are a common choice for ADS on today’s blockchains [21]. In a database of N items, updating a single state entry in an MPT has a time complexity of $O(\log(N))$ [15]. However, MPT and other existing trie-based ADSes suffer from large proofs and a dependency on the validator node having a large amount of physical memory to avoid excessively reading from the SSD. At the same time, MPTs are not suitable for storage on flash storage, as the randomly distributed update-heavy workload results in high WA. To top it off, the worst-case size for inclusion and exclusion proofs can be quite large. These factors result in Merkleization becoming a significant bottleneck that limits the overall throughput of the execution layer and the blockchain.
AVL tree based ADSes are popular alternatives to MPTs, as they achieve faster updates, lookups, and proof generation due to the self-balancing AVL tree. AVL trees provide a marginal performance increase over MPTs in the average case, but still suffer from O(log N) tree nodes modifications per state update.
@@
NOMT is a state-of-the-art ADS that uses a flash-optimized layout for a binary Merkle tree with compressed metadata, overcoming some limitations of existing MPT-based ADS implementations. This design results in better performance than existing solutions and has garnered interest in the space. NOMT implements an array of improvements including tree arity, flash native layout, a write-ahead log, and caching. However, it continues to face similar asymptotic limitations and write amplification as AVL tree-based ADSes.
@@
Table 1: Fields in a QMDB entry. ID and Version are 8 bytes, Key is 32 bytes, and Value can hold up to $2^{24}$ bytes.
@@ 3 QMDB Design
QMDB is architected as a binary Merkle tree illustrated in Figure 1. At the top is a single global root that connects a set of shard roots, each of which represents the subtree of the world state that is managed by an independent QMDB shard. The shard root itself is connected to a set of upper nodes, which, in turn, are connected to fixed-size subtrees called twigs; each of these twigs has a root that stores the Merkle hash of the subtree and a bitmap called ActiveBits to track which entries are part of the most current world state. Entries (the twig’s leaves) are immutable, making it unnecessary to read or write the entry root during Merkleization; this results in QMDB only ever reading/writing the global root, shard roots, upper nodes, and twig roots during Merkleization. The twig essentially compresses the actual state keys and values into a single hash and bitmap, making the data required for Merkleization small enough to fit in a small amount of DRAM rather than being stored on SSD.
@@ 3.1 Entries and Twigs
Twigs are subtrees within QMDB’s Merkle Tree; each twig has a fixed depth, by extension a fixed number of entries stored in the leaf nodes (2048 in our implementation). A set of upper nodes connects all twigs to a single shard root, with null nodes to represent uninitialized values; these upper nodes are immutable once all their descendant entries have been initialized. In addition to the actual Merkle subtree, Twigs also store the Merkle hash of their root node and ActiveBits, a bitmap that describes whether each contained entry contains state that has not been overwritten or deleted. The twig essentially compresses the information required to Merkleize 2048 entries and their upper nodes ($\geq 256kb$) into a single 32-byte hash and a 256-byte bitmap (99.9% compression). This compression is the key to enabling fully in-memory Merkleization in QMDB.
@@
Each twig follows a lifecycle of 4 states: Fresh, Full, Inactive, and Pruned (Table 2). An example of the layout of QMDB’s state tree is presented in Figure 1 There is exactly one fresh twig per shard, and entries are always appended to the fresh twig. After all entries in the twig are marked inactive as a result of update and delete operations, the twig transitions into the inactive state before eventually being pruned and replaced by the Merkle hash of the root. Upper nodes that contain only pruned twigs are recursively pruned, further reducing the memory footprint of QMDB.
The grouping of entries into twigs reduces the DRAM footprint of QMDB to the degree that all nodes affected by Merkleization can be stored in a small amount of DRAM. In a hypothetical scenario with $2^{30}$ entries (approx. 1 billion), the system must keep at most $2^{19}$ 288-byte (32-byte twig root hash & 2048-bit ActiveBits bitmap) full twigs, 1 fresh twig and $12^{19}-$ 1 32-byte (node hash) upper nodes totaling around 160 megabytes. In practice, the majority of the $2^{19}$ twigs will be pruned, resulting in the average size being much smaller.
@@
QMDB stores an entry every time state is modified, making the state tree grow proportionally to the number of state modifications. To combat this tree growth, a dedicated compressor worker periodically compresses QMDB’s state tree by removing and re-appending old entries to the fresh twig, accelerating the progression of the twig lifecycle and allowing more subtrees to be pruned. The compression is deterministic and is triggered upon updates and insertions.
QMDB’s Merkle proof size and proof generation complexity grow proportionally to $(\log(U))^{2}$ of the number of state updates ($U$) rather than the number of unique keys ($K$) due to its append-only nature. However, the ratio of $U$ to $K$ remains small enough that the order-of-magnitude improvement in Merkleization performance dominates the small additional cost. In addition, ZK-proofs can be used to compress the proof witness data which drastically reduces proof verification cost, avoiding end-to-end bottlenecks in the proof size.
@@ 3.2 Indexer
The indexer maps the application-level keys to their respective entries, enabling QMDB’s CRUD interface. To support efficient insertion and deletion of entries (Section 3.3), the indexer must support ordered key iteration. The indexer can be freely swapped for different implementations depending on specific application needs, but we expect that QMDB’s default in-memory indexer will meet the resource requirements of the majority of use cases.
QMDB’s default indexer consumes about 15.4 bytes of DRAM per key and serves key lookups in-memory to reduce SSD I/Os; this is achieved by keeping only the 9 most significant bytes of the key, allowing a slightly higher probability of key collision and trading worst-case performance for reduced DRAM usage. Using just 8 gigabytes of DRAM, the in-memory indexer can index more than 1 billion entries, making it suitable for a wide range of applications. We chose the B-tree map as the basis for the underlying structure of QMDB’s default indexer to take advantage of B-tree’s high cache locality, low memory overhead, support for ordered key iteration, and graceful handling of key collisions. We use fine-grained reader-writer locks (determined by the first two bytes of the key hash) to minimize contention when updating entries.
@@ 3.3 CRUD interface
QMDB exposes a CRUD (Create, Read, Update, Delete) interface, and in this section we provide a high-level overview of how each operation is implemented. In all examples, we present the operation of the system when using the default in-memory indexer; other indexers may require more reads or writes to serve the same workload. For each operation, we present an intuitive explanation followed by a more formal description along with a description of the SSD I/O required to synchronously handle the request. All writes in QMDB are buffered in twigs (DRAM) and persisted to SSD in batches, so each SSD write is amortized across 2048 entries; to precisely express the cost of each operation, we refer to a entry write as $\frac{1}{2048}$ of a single batched flush to SSD. For brevity, we omit the Id, Version, and Value fields when describing new entries, so an entry $E$ is defined as:
@@ 3.5 Parallelization
QMDB shards its key space into contiguous spans with sentry nodes to define logical boundaries that prevent state modifications from crossing shard boundaries (i.e., PrevKey and NextKey will always fall within the same shard). This sharding enables QMDB to better saturate underlying hardware resources and scale to bigger or multiple physical servers.
@@ 4 Evaluation
In this section, we present a preliminary evaluation of the performance of QMDB and compare it to RocksDB and NOMT. On a comparable workload and evaluation setup, QMDB achieves 6×6$\times$6 × higher updates per second than RocksDB and 8×8$\times$8 × higher updates per second than NOMT.
@@
Figure 3 shows the throughput of QMDB compared to RocksDB, demonstrating that QMDB delivers 6×6$\times$6 × more updates per second than RocksDB. This speedup is in fact an underestimate of QMDB’s advantage over two-layer systems, given that all benchmarks compare QMDB with Merkleization to RocksDB without Merkleization. We believe the primary factor driving this speedup to be QMDB’s single-layer design with no in-place overwriting of persistent data. This design trades off functionality unnecessary for blockchain workloads for extra throughput. Examples of features and characteristics of RocksDB that are not required in blockchain workloads include efficient range/prefix queries and spatial locality of key-value pairs.
@@
Figure 3: QMDB shows a 6×6$\times$6 × increase in throughput over RocksDB. QMDB is able to do 601K updates/sec with 6 billion entries and demonstrates superior performance across all operation types. These results were obtained on an AWS instance with 2 SSDs and 64 vCPUs.
@@ 4.2 Up to 8X throughput vs state-of-the-art
We note that QMDB’s performance relies on its indexer, which incurs some DRAM overhead. Compared to NOMT’s overhead of 1 byte per entry, QMDB incurs an additional 14 bytes per entry with its in-memory indexer and an additional 1–2 bytes per entry with its hybrid indexer). We consider this to be a reasonable trade-off given the 8×8$\times$8 × increase in throughput, with QMDB’s hybrid indexer still offering a speedup for DRAM-constrained setups.
@@ 4.4 Scaling up and down
Scaling up to hundreds of billions of entries. The hybrid indexer trades off SSD capacity and system throughput to reduce the DRAM footprint of the QMDB indexing layer to just 2–3 bytes per entry, allowing servers with a high ratio of SSD capacity to DRAM capacity to scale to huge world states. Table LABEL:table:eval:aws-datasize shows the maximum theoretical number of entries that can be stored in QMDB running on various different AWS instances. We calculate that the i3en.metal instance with high SSD capacity and a reasonable amount of DRAM could scale to 280 billion entries, far exceeding the needs of any existing production blockchain. Due to the prohibitive amount of time necessary to populate hundreds or even tens of billions of keys, we only run experiments up to 15 billion entries and conservatively extrapolate the results. The average DRAM overhead actually drops as more entries are inserted; 1 billion entries cost about 3 bytes of DRAM per entry, which drops to just 2.2 bytes per entry for 15 billion entries, indicating that the marginal DRAM overhead per additional entry is close to constant.

### v2

@@
Abstract
Quick Merkle Database (QMDB) addresses longstanding bottlenecks in blockchain state management by integrating key-value (KV) and Merkle tree storage into a single unified architecture. QMDB delivers a significant throughput improvement over existing architectures, achieving up to 6×6$\times$6 × over the widely used RocksDB and 8×8$\times$8 × over NOMT, a leading verifiable database. Its novel append-only twig-based design enables one SSD read per state access, $O(1)$ IOs for updates, and in-memory Merkleization on a memory footprint as small as 2.3 bytes per entry enabling it to run on even modest consumer-grade PCs. QMDB scales seamlessly across both commodity and enterprise hardware, achieving up to 2.28 million state updates per second. This performance enables support for 1 million token transfers per second (TPS), marking QMDB as the first solution achieving such a milestone. QMDB has been benchmarked with workloads exceeding 15 billion entries (10× Ethereum’s 2024 state) and has proven the capacity to scale to 280 billion entries on a single server. Furthermore, QMDB introduces historical proofs, unlocking the ability to query its blockchain’s historical state at the latest block. QMDB not only meets the demands of current blockchains but also provides a robust foundation for building scalable, efficient, and verifiable decentralized applications across diverse use cases.
@@ 1 Introduction
Updating, managing, and proving world state are key bottlenecks facing the execution layer in modern blockchains. Within the execution layer, the storage layer, in particular, has traditionally traded off performance (throughput) and decentralization (capital and infrastructure barriers to participation). Blockchains typically implement state management using an Authenticated Data Structure (ADS) such as a Merkle Patricia Trie (MPT). Unfortunately, typical MPT-based ADSes incur a high amount of write amplification (WA) with many costly random writes for each state update, which requires storing the entire structure in DRAM to avoid getting bottlenecked by the SSD. As a result, the performance and scaling of blockchains is I/O-bound, and the key to unlocking higher performance with larger datasets is to optimize the use of SSD IOPS more efficiently and reduce WA.
We present Quick Merkle Database (QMDB), a resource-efficient SSD-optimized ADS with in-memory Merkleization that implements a superset of the app-level features of existing RocksDB-backed MPT ADSes with 6×6$\times$6 × throughput on large datasets. Qmdb performs state reads with a single SSD read, state updates with O(1) IO, and performs Merkleization fully in-memory with no SSD reads or writes. These operations are theoretically optimal regarding disk IO complexity. Additionally, QMDB has a DRAM footprint small enough to run on consumer-grade PCs.
@@
QMDB eliminates this inefficiency by unifying the world state and Merkle tree storage, persisting all state updates in an append-only log, and eliminating all SSD reads and writes from Merkleization. By grouping updates into fixed-size immutable subtrees called twigs, QMDB can Merkleize state updates without reading or writing any world state; this essentially compresses the Merkle tree by several orders of magnitude, allowing it to be stored in a modest amount of DRAM. QMDB leverages typical blockchain workload characteristics to eliminate features commonly found in KVDBs—such as key iterations—thereby reducing performance bottlenecks.
@@
Figure 1: Entries are inserted sequentially into the leaves of the Fresh twig, and all leaves have the same depth. The twig eventually transitions into the Full state. As Entries are deleted, Full twigs become Inactive, then transition to Pruned. Upper nodes are recursively pruned after both of their children are pruned.
@@ 2 Background
MPTs combine the efficient proof generation of the Merkle tree with the fast lookups of the Patricia trie and are a common choice for ADS on today’s blockchains [23]. In a database of N items, updating a single state entry in an MPT has a time complexity of $O(\log(N))$ [17]. However, MPT and other existing trie-based ADSes suffer from large proofs and a dependency on the client having a large amount of physical memory to avoid excessive random SSD reads. At the same time, MPTs are not suitable for storage on flash storage, as the randomly distributed update-heavy workload results in high WA. To top it off, the worst-case size for inclusion and exclusion proofs can be quite large. These factors result in Merkleization becoming a significant bottleneck that limits the overall throughput of the execution layer and the blockchain.
AVL tree based ADSes are popular alternatives to MPTs, as they achieve faster updates, lookups, and proof generation due to the self-balancing AVL tree. The AVL tree is path-dependent, unlike the MPT, meaning its state root is influenced by the specific sequence of state change actions. AVL trees provide a marginal performance increase over MPTs in the average case, but still suffer from O(log N) tree nodes modifications per state update.
LVMT[13] proposes a layered storage model to reduce the space and complexity of maintaining authenticated blockchain states. By partitioning the state into multiple segments and using cryptographic accumulators, it compresses less frequently accessed data while preserving verifiability. Proof generation becomes simpler, as intermediate accumulators shorten authentication paths. However, integrating multiple layers increases system complexity and demands careful configuration—suboptimal settings can lead to poor performance. Furthermore, LVMT depends on well-optimized cryptographic primitives.
@@
NOMT is a state-of-the-art ADS that uses a flash-optimized layout for a binary Merkle tree with compressed metadata, overcoming some limitations of existing MPT-based ADS implementations. NOMT implements an array of improvements including tree arity, flash native layout, a write-ahead log, and caching. This design results in better performance than existing solutions and has garnered interest in the space. However, NOMT remains an implementation-level optimization of MPT, offering only constant-factor reductions in disk I/O. It still faces inherent asymptotic limitations and write amplification issues. Additionally, it is affected by the key sparsity problem commonly observed in trie-based structures.
@@
Table 1: Fields in a QMDB entry. ID and Version are 8 bytes. Key has up to $2^{8}$ bytes and Value can hold up to $2^{24}$ bytes.
@@ 3 QMDB Design
QMDB is architected as a binary Merkle tree illustrated in Figure 1. At the top is a single global root that connects a set of shard roots, each of which represents the subtree of the world state that is managed by an independent QMDB shard. The shard root itself is connected to a set of upper nodes, which, in turn, are connected to fixed-size subtrees called twigs; each of these twigs has a root that stores the Merkle hash of the subtree and a bitmap called ActiveBits to track which entries are part of the most current world state. The twig root is determined by the sequence of entries, making it path-dependent. Entries (the twig’s leaves) are append-only and immutable, making it unnecessary to read or write the entry root during Merkleization; this results in QMDB only ever reading/writing the global root, shard roots, upper nodes, and twig roots during Merkleization. The twig essentially compresses the actual state keys and values into a single hash and bitmap, making the data required for Merkleization small enough to fit in a small amount of DRAM rather than being stored on SSD.
@@ 3.1 Entries and Twigs
Twigs are subtrees within QMDB’s Merkle Tree; each twig has a fixed depth, by extension a fixed number of entries stored in the leaf nodes of the same depth (2048 in our implementation). A set of upper nodes connects all twigs to a single shard root, with null nodes to represent uninitialized values; these upper nodes are immutable once all their descendant entries have been initialized. In addition to the actual Merkle subtree, Twigs also store the Merkle hash of their root node and ActiveBits, a bitmap that describes whether each contained entry contains state that has not been overwritten or deleted. The twig essentially compresses the information required to Merkleize 2048 entries and their upper nodes ($\geq 256kb$) into a single 32-byte hash and a 256-byte bitmap (99.9% compression). This compression is the key to enabling fully in-memory Merkleization in QMDB.
@@
Each twig follows a lifecycle of 4 states: Fresh, Full, Inactive, and Pruned (Table 2). An example of the layout of QMDB’s state tree is presented in Figure 1 There is exactly one fresh twig per shard, and entries are always appended to the fresh twig. After all entries in the twig are marked inactive as a result of update and delete operations, the twig transitions into the inactive state before eventually being pruned and replaced by the Merkle hash of the root. Upper nodes that contain only pruned twigs are recursively pruned, further reducing the memory footprint of QMDB; a dedicated garbage collection thread duplicates old valid entries into the fresh twig, reducing fragmentation and allowing larger subtrees to be pruned. In theory, once the entire subtree originating at a child of the shard root is pruned, the root itself can be pruned to reduce the depth of the tree by one.
The grouping of entries into twigs reduces the DRAM footprint of QMDB to the degree that all nodes affected by Merkleization can be stored in a small amount of DRAM. In a hypothetical scenario with $2^{30}$ entries (approx. 1 billion), the system must keep at most $2^{19}$ ($\frac{2^{30}}{2048}$) 288-byte (32-byte twig root hash & 2048-bit ActiveBits bitmap) full twigs, 1 fresh twig and $12^{19}-$ 1 32-byte (node hash) upper nodes totaling around 160 megabytes. In practice, the majority of the $2^{19}$ twigs will be pruned, resulting in the average size being much smaller.
@@
QMDB stores an entry every time state is modified, making the state tree grow proportionally to the number of state modifications. To combat this tree growth, a dedicated compaction worker periodically compacts QMDB’s state tree by removing and re-appending old entries to the fresh twig, accelerating the progression of the twig lifecycle and allowing more subtrees to be pruned. The compaction logic must be deterministic when used in a consensus system or for stateless validation. The current implementation ensures that the active entry ratio per shard remains above a predefined threshold, triggering compression during updates and insertions.
QMDB’s Merkle proof size and proof generation complexity grow proportionally to $\log(U)$ of the number of state updates ($U$) rather than the number of unique keys ($K$) due to its append-only nature. However, the ratio of $U$ to $K$ remains small enough that the order-of-magnitude improvement in Merkleization performance dominates the small additional cost. Assuming 10,000 transactions per second and an average of 5 KV updates per transaction, the tree depth after one year will be at most 41 ($log2(10000*5*3600*24*365)$); however, in practice the actual depth will be much shallower due to pruning of overwritten subtrees and garbage collection. In addition, ZK-proofs can be used to compress the proof witness data which drastically reduces proof verification cost, avoiding end-to-end bottlenecks in the proof size.
@@ 3.2 Indexer
The indexer maps the application-level keys to their respective entries, enabling QMDB’s CRUD interface. To support efficient insertion and deletion of entries (Section 3.3), the indexer must support ordered key iteration. The indexer can be freely swapped for different implementations depending on specific application needs, but we expect that QMDB’s default in-memory indexer will meet the resource requirements of the majority of use cases. This modularity potentially enables optimizations to increase the performance or memory efficiency of the indexer such as those found in systems such as SILT [15] or MICA [16].
QMDB’s default indexer consumes approximately 15.4 bytes of DRAM per key and serves key lookups in-memory to minimize SSD I/Os. This efficiency is achieved by using only the 9 most significant bytes of each key, which slightly increases the likelihood of key collisions but strategically trades worst-case performance for reduced DRAM usage. Of these 9 bytes, the first 2 bytes serve as the sharding key for the indexer, leaving a 7-byte memory footprint for key storage. The remaining 8.4 bytes consist of a 6-byte SSD position offset and additional data structure overhead, which is amortized across all keys. Using just 16 gigabytes of DRAM, the in-memory indexer can index more than 1 billion entries, making it suitable for a wide range of applications. We chose the B-tree map as the basis for the underlying structure of QMDB’s default indexer to take advantage of B-tree’s high cache locality, low memory overhead, support for ordered key iteration, and graceful handling of key collisions. We use fine-grained reader-writer locks (determined by the first two bytes of the key hash) to minimize contention when updating entries.
@@ 3.3 CRUD interface
QMDB exposes a CRUD (Create, Read, Update, Delete) interface, and in this section we provide a high-level overview of how each operation is implemented. In all examples, we present the operation of the system when using the default in-memory indexer; other indexers may require more reads or writes to serve the same workload. For each operation, we present an intuitive explanation followed by a more formal description along with a description of the SSD I/O required to synchronously handle the request. All writes in QMDB are buffered in twigs (DRAM) and persisted to SSD in batches, so each SSD write is amortized across 2048 entries; to precisely express the cost of each operation, we refer to a entry write as $\frac{1}{2048}$ of a single batched flush to SSD. For brevity, we omit the Id, Version, and Value fields when describing new entries (see Table 1), so an entry $E$ is defined as:
@@ 3.5 Parallelization
QMDB shards its key space into contiguous spans using the most significant bits—for example, the first 4 bits can create 16 shards—with boundary nodes to define logical boundaries that prevent state modifications from crossing shard boundaries (i.e., PrevKey and NextKey will always fall within the same shard). This sharding enables QMDB to better saturate underlying hardware resources and scale to bigger or multiple physical servers.
@@ 4 Evaluation
In this section, we present a preliminary evaluation of the performance of QMDB and compare it to RocksDB and NOMT. On a comparable workload and evaluation setup, QMDB achieves 6×6$\times$6 × higher updates per second than RocksDB and 8×8$\times$8 × higher updates per second than NOMT. When measuring the performance of QMDB, we generate 100,000 transactions per block–each creating 10 entries–and run the workload for 7000 blocks to create a total of 7 billion entries. Periodically (every billion entries) we test the throughput and latency of reads, updates, deletions, and creations, and after all 7 billion entries are populated we measure transactions per second (TPS) using transactions consisting of 9 writes, 15 reads, 1 create, and 1 delete.
@@
Figure 3 shows the throughput of QMDB compared to RocksDB (storing the application-level key-values with no Merkleization), demonstrating that QMDB delivers 6×6$\times$6 × more updates per second than RocksDB. This speedup is in fact an underestimate of QMDB’s advantage over RocksDB-based systems, given that all benchmarks compare QMDB with Merkleization to RocksDB without Merkleization. We believe the primary factor driving this speedup to be QMDB trading off functionality unnecessary for blockchain workloads for extra throughput. Examples of features and characteristics of RocksDB that are not required in blockchain workloads include efficient range/prefix queries and spatial locality of key-value pairs.
@@
Figure 3: QMDB shows a 6×6$\times$6 × increase in throughput over RocksDB. QMDB is able to do 601K updates/sec with 6 billion entries and demonstrates superior performance across all operation types. These results were obtained on an AWS c7gd.metal instance with 2 SSDs and 64 vCPUs.
@@ 4.2 Up to 8X throughput vs state-of-the-art
We note that QMDB’s performance relies on its indexer, which incurs some DRAM overhead. Compared to NOMT’s overhead of 1–2 bytes per entry, QMDB incurs an additional 14 bytes per entry with its in-memory indexer and an additional 1–2 bytes per entry with its hybrid indexer). We consider this to be a reasonable trade-off given the 8×8$\times$8 × increase in throughput, with QMDB’s hybrid indexer still offering a speedup for DRAM-constrained setups.
@@ 4.4 Scaling up and down
Scaling up to hundreds of billions of entries. The hybrid indexer trades off SSD capacity and system throughput to reduce the DRAM footprint of the QMDB indexing layer to just 2.3 bytes per entry, allowing servers with a high ratio of SSD capacity to DRAM capacity to scale to huge world states. Table LABEL:table:eval:aws-datasize shows the maximum theoretical number of entries that can be stored in QMDB running on various different AWS instances. We calculate that the i3en.metal instance with high SSD capacity and a reasonable amount of DRAM could scale to 280 billion entries, far exceeding the needs of any existing production blockchain. Due to the prohibitive amount of time necessary to populate hundreds or even tens of billions of keys, we only run experiments up to 15 billion entries and conservatively extrapolate the results. The average DRAM overhead actually drops as more entries are inserted; 1 billion entries cost about 3 bytes of DRAM per entry, which drops to just 2.2 bytes per entry for 15 billion entries, indicating that the marginal DRAM overhead per additional entry is close to constant.